OpenStudy (anonymous):

Given cost and revenue functions, find the profit function. P(x) = R(x)-C(x)

OpenStudy (anonymous):

\[R(x)=\frac{ 575x }{ \sqrt{x} }-3x\]

OpenStudy (anonymous):

\[C(x)=2000+150x-0.002x^2\]